# Kai Shu, Lu Cheng, Huan Liu

### v3

@@ 5.1.3 Emerging Tasks
Recently, we have seen more customized utilization of LLM-as-a-judge to evaluate emerging tasks such as code understanding Zhao et al. (2024a), legal knowledge Fei et al. (2023), game development Isaza-Giraldo et al. (2024), ocean science Bi et al. (2023), healthcare conversations Wang et al. (2024j), debating judgment Liang et al. (2024a), retrival-augmented generation Dhole et al. (2024), cross ability Zhong et al. (2024) and more. This trend reflects the growing adaptability of LLMs-as-judges in evaluating diverse and specialized domains.
@@ 5.2.1 Larger Models as Judges
One intuitive idea for adopting LLM-as-a-judge in alignment tuning is to leverage the feedback from larger, more powerful LLMs to guide smaller, less capable models. Bai et al. (2022) first propose employing the AI’s feedback to build a harmless AI assistant. They train the reward model using synthetic preference data based on the preference of a pre-trained language model. Building on this, Lee et al. (2023) discover that the RLAIF method can achieve comparable performance with RLHF even when the LLM judge is not strong enough. They also introduced DIRECT-RLAIF, which directly employs an off-the-shelf LLM as the judge model to mitigate reward staleness in reward models. To avoid reward hacking in alignment, Sun et al. (2024) devise an instructable reward model trained on synthetic preference data. It enables humans to perform RL-time interventions to better align the target policy with human values. Apart from the abovementioned studies, Guo et al. (2024) introduce online AI feedback (OAIF), directly utilizing the preference signals from an annotation model to train the target model. There are also works that utilize multi-agent cooperation for better judgment in alignment tuning. Arif et al. (2024) and Sengupta et al. (2024) construct a synthetic preference optimization dataset using multi-agent workflows and adopt LLMs as judges with diverse prompting strategies and pipelines. Similarly, Li et al. (2024i) employ multiple LLMs to debate with each other, iteratively improving response quality, while creating a judge LLM to select preferred responses for enhanced instruction tuning. To align the generated code with human preference, Weyssow et al. (2024) introduce CodeUltraFeedback, a preference coding dataset constructed using the LLM-as-a-Judge methodology. This synthetic dataset is later used to fine-tune and align small code LLMs using SFT and DPO. Recently, Wang et al. (2024d) propose BPO, using GPT-4 as the judge and constructing synthetic pairwise feedback for knowledge depth and breadth balance in the alignment process.
@@ 5.2.2 Self-Judging
Another line of work aims to utilize the preference signal from the same LLM for self-improving. Yuan et al. (2024c) first propose the concept of a self-rewarding LLM, where pairwise data is constructed by having the LLM itself serve as the judge. Following them, Wu et al. (2024a) introduce meta-rewarding that judges the LLMs’ judge and uses the feedback to refine their judging skills. Their LLM-as-a-meta-judge approach significantly enhances the model’s capacity to evaluate and follow instructions. To improve synthetic data quality, Pace et al. (2024) combine the Best-of-N and Worst-of-N sampling strategies and introduce the West-of-N approach. Lee et al. (2024) devise Judge augmented Supervised Fine-Tuning (JSFT) to train a single model to act as both a policy and a judge. To fully utilize this judge model, they also propose the Self-Rejection by Tournament method to select the best response in the inference time. Unlike the above approaches, which use the LLM as a judge for pairwise data construction, Tong et al. (2024) apply LLM-as-a-judge as a self-filtering method to ensure the quality of synthetic data pairs in alignment tasks for reasoning. To reduce the computation overhead in pairwise judgment, Zhai et al. (2024) propose a ranked pairing method for self-preferring language models, which accelerates the comparison process by measuring the strength of each response against a baseline. Liu et al. (2024c) introduce meta-ranking, enabling weaker LLMs to serve as reliable judges and provide trustworthy feedback. They also apply their meta-ranking method in post-SFT training, combining it with Kahneman-Tversky Optimization (KTO) for improved alignment. To enhance the quality of the synthetic instruction tuning data, Liang et al. (2024c) introduce an iterative self-enhancement paradigm (I-SHEEP). During training, they adopt LLM-as-a-judge to score the synthetic responses and set a threshold to collect high-quality query-response pairs for the subsequent training iteration. Recently, Yasunaga et al. (2024) propose to combine LLM-as-a-judge and data synthesis and successfully build human-aligned LLMs with a few annotation.
Several works have also employed LLM-as-a-judge in specific domains or for particular attributes. Zhang et al. (2024g) propose a self-evaluation mechanism, judging responses’ factuality by generating question-answer pairs. They then utilize these self-annotated responses to fine-tune the model via the DPO algorithm for better factuality. In robotics, Zeng et al. (2024) iteratively update the reward function with the self-ranking responses from LLMs, boosting learning efficiency without human supervision. In the multimodal domain, Ahn et al. (2024) propose iterative self-retrospective judgment (i-SRT), which employs self-reflection to improve both response generation and preference modeling.
@@ 5.3.1 Traditional Retrieval
Beyond general retrieval tasks, LLMs have demonstrated their utility as judges in specialized applications. For example, Ma et al. (2024) outline a few-shot workflow that employs a general-purpose LLM for relevance judgments in legal information retrieval. This model achieves high consistency with expert annotations by breaking down tasks into stages, facilitating the integration of expert reasoning to improve relevance assessment accuracy in legal contexts.
@@ 5.3.2 Retrieval-Augmented Generation (RAG)
RAG systems tailored to specific domains reveal the potential for LLMs to navigate complex queries by integrating specialized knowledge structures. In domain-specific retrieval, Wang et al. (2024b) present BIORAG, an advanced RAG framework that enhances vector retrieval with hierarchical knowledge structures. BIORAG adopts a self-aware evaluated retriever to continuously judge the adequacy and relevance of the information it has collected, thus improving the accuracy of the retrieved document. For biomedical research, Li et al. (2024c) introduce DALK (Dynamic Co-Augmentation of LLMs and Knowledge Graphs), a novel system that combines an LLM with a continuously evolving Alzheimer’s Disease (AD) knowledge graph derived from the scientific literature. Using the novel self-aware knowledge retrieval method, DALK employs the judging capability of LLMs to do noise filtering, enhancing the LLM’s inferential performance in AD-related queries. Similarly, Jeong et al. (2024) propose Self-BioRAG, a framework that adapts RAG principles to biomedical applications. Self-BioRAG employs the LLM to select the best evidence and generate the answer based on the selected evidence and encoded knowledge.
@@ 5.4.1 Reasoning Path Selection
Some works focus on sample-level selection in the reasoning process.  Kawabata and Sugawara (2024) introduce REPS (Rationale Enhancement through Pairwise Selection), judging and selecting valid rationales via pairwise self-evaluation with LLMs and training a verifier based on these data. Another reasoning path selection problem for LLMs is diversity, Lahoti et al. (2023) find that LLMs grasp the concept of diversity and can identify aspects where responses lack diversity. By selecting and aggregating multiple critiques, LLMs can achieve similar gains, compared to multiple iterations of critique and revision. In multi-agent collaboration frameworks, Liang et al. (2023) propose multi-agent debating (MAD), a novel paradigm to facilitate debating and discussion among multiple agents. They leverage the judge LLM to select the most reasonable response as the final output at the end of the debating process. Similarly, Li et al. (2024b) propose new roles in layer-based multi-agent collaboration, employing the judge LLM to select the high-quality and reasonable response and thus enhancing the whole system’s token utilizing efficiency by a considerable margin.
@@
Additionally, there are also many works that focus on step-level reasoning path selection, leveraging the judge LLM as the process reward model (PRM) to evaluate state scores. Creswell et al. (2023) decompose the reasoning process into Selection and Inference. In the selection step, they leverage the LLM itself to judge and evaluate each potential reasoning trace, selecting the appropriate one for the following inference step. Xie et al. (2024b) propose the Kwai-STaR framework, which transforms LLMs into state-transition reasoners to judge and select the best reasoning state for themselves in mathematical reasoning. Lightman et al. (2023) train a LLM as the PRM to conduct inference-time supervision and perform best-of-N sampling strategy during reasoning stage. Following them, Setlur et al. (2024) further propose process advantage verifiers (PAVs), generating rewards based on the changes in the likelihood of producing correct responses in the future. Other works simulate advanced cognitive structures as the reasoning process. Hao et al. (2023) employ LLMs as a world model to simulate the state of the environment and perform Monte Carlo Tree Search (MCTS) to improve performance on tasks requiring deliberate path selection. Furthermore, Besta et al. (2024) takes the outputs generated by LLMs as an arbitrary graph. An LLM thought is modeled as a vertex, while an edge is a dependency between thoughts. This framework enables a systematic judge of coherence and logical reasoning for each reasoning state. Yao et al. (2023a) propose Tree-of-Thoughts (ToT), where each thought serves as an intermediate step toward problem-solving. It decomposes reasoning into steps, self-evaluates and judges progress at each state, and uses search algorithms with LMs to judge thought paths through lookahead and backtracking.
@@
EVALBIAS-BENCH Wang et al. (2024c) Alignment 80 Human Accuracy Various Bias
@@
Arena-Hard Auto Li et al. (2024j) Alignment 500 GPT-4-Turbo Separability, Agreement Challenging
@@ 5.4.2 Reasoning with External Tools
Yao et al. (2023b) first propose to use LLMs in an interleaved manner to generate reasoning traces and task-specific actions. Reasoning traces help the model to judge and update action plans, while actions enable it to interact with external sources. Further, Auto-GPT was introduced by Yang et al. (2023) to deliver more accurate information with LLM-as-a-judge for tool-using. By equipping with a range of external complex tools, LLMs become more versatile and capable, improving planning performance by judging and reasoning which tools to use. Sha et al. (2023) explore LLMs’ potential in reasoning and judging, employing them as decision-making components for complex autonomous driving scenarios that require human commonsense understanding. Zhou et al. (2024b) utilize a self-discovery process where LLMs perform judgment based on the given query and select the most feasible reasoning structure for the following inference stage.
@@ 6.1 General Performance
  • JUDGE-BENCH Tan et al. (2024a), DHP Wang et al. (2024g), RewardBench Lambert et al. (2024) and SOS-BENCH Penfever et al. (2024), which operate at larger scales, utilizing metrics like Cohen’s kappa, Discernment Score, and normalized accuracy to benchmark general LLM performance.
@@ 6.2 Bias Quantification
Mitigating bias in LLM judgments is critical to ensuring fairness and reliability. Typical benchmarks include EVALBIAS-BENCH Wang et al. (2024c) and CALM Ye et al. (2024a), focus explicitly on quantifying biases, including those emerging from alignment and robustness under adversarial conditions. Besides, Shi et al. (2024a) evaluate metrics such as position bias and percent agreement in question-answering tasks.
@@ 6.3 Challenging Task Performance
Benchmarks designed for difficult tasks push the boundaries of LLM evaluation. For example, Arena-Hard Auto Li et al. (2024j) and JudgeBench Tan et al. (2024a) select harder questions based on LLMs’ performance for conversational QA and various reasoning tasks, respectively. CALM Ye et al. (2024a) explores alignment and challenging scenarios, using metrics such as separability, agreement, and hacked accuracy to evaluate performance in manually identified hard data sets.
@@ 6.4 Domain-Specific Performance
Domain-specific benchmarks provide task-focused evaluations to assess LLMs’ effectiveness in specialized contexts. Concretely, Raju et al. (2024) measure separability and agreement across tasks, leveraging metrics such as Brier scores for insights in specific domains such as coding, medical, finance, law and mathematics. CodeJudge-Eval Zhao et al. (2024a) specifically evaluates LLMs for judging code generation with execution-focused metrics such as accuracy and F1 score. This idea has also been adopted by several following works in code summarization and generation evaluation Wu et al. (2024b); Yang et al. (2024); Tong and Zhang (2024)
@@ 7.1 Bias & Vulnerability
In addition to these general biases, specific evaluative biases emerge when LLMs act as judges. Order Bias is a prominent issue where the sequence of candidates influences preferences Zheng et al. (2023); Wang et al. (2023c); Koo et al. (2023a); Shi et al. (2024b). This bias can distort evaluation outcomes, particularly in pairwise comparisons, and is more pronounced when the quality gap between competing responses is small Shi et al. (2024b); Wang et al. (2023c). Egocentric Bias arises when LLMs favor outputs generated by the same model, compromising objectivity Liu et al. (2023c); Koo et al. (2023a); Wataoka et al. (2024). This issue is particularly pronounced when evaluation metrics are designed using the same model, leading to inflated scores for outputs from the originating model Liu et al. (2023c). Length Bias, another prevalent challenge, skews evaluations by disproportionately favoring longer or shorter responses regardless of quality Zheng et al. (2023); Koo et al. (2023a). Additional biases, such as Misinformation Oversight Bias, Authority Bias, and Beauty Bias, further complicate LLM evaluations. For instance, misinformation oversight bias reflects a tendency to overlook factual errors, authority bias favors statements from perceived authoritative sources, and beauty bias prioritizes visually appealing content over substantive quality Chen et al. (2024b, d); Stephan et al. (2024). Similarly, Verbosity Bias shows a preference for lengthier explanations, often equating verbosity with quality, which can mislead the judgment process Yuan et al. (2024a). Furthermore, Sentiment Bias skews evaluations based on emotional tone, favoring responses with positive phrasing Ye et al. (2024a).
LLM judges are also highly susceptible to adversarial manipulations. Techniques like JudgeDeceiver highlight the risks posed by optimization-based prompt injection attacks, where carefully crafted adversarial sequences can manipulate LLM judgments to favor specific responses Shi et al. (2024a); BENCHMARK ; Banerjee et al. (2024). Similarly, universal adversarial phrases can drastically inflate scores in absolute scoring paradigms, revealing vulnerabilities in zero-shot assessment setups Liusie et al. (2023); Raina et al. (2024); Doddapaneni et al. (2024). These manipulations raise concerns about the reliability of LLM judges in high-stakes scenarios such as leaderboards and academic assessments Shi et al. (2024a); Raina et al. (2024).

### v4

@@ 5.1.3 Emerging Tasks
Recently, we have seen more customized utilization of LLM-as-a-judge to evaluate emerging tasks such as code understanding Zhao et al. (2024a), legal knowledge Fei et al. (2023), game development Isaza-Giraldo et al. (2024), ocean science Bi et al. (2023), healthcare conversations Wang et al. (2024j), debating judgment Liang et al. (2024a), retrival-augmented generation Dhole et al. (2024), cross ability Zhong et al. (2024), HCI Li et al. (2024j), role-playing Zhou et al. (2024b), RAG Jin et al. (2024) and more. This trend reflects the growing adaptability of LLMs-as-judges in evaluating diverse and specialized domains.
@@ 5.2.1 Larger Models as Judges
One intuitive idea for adopting LLM-as-a-judge in alignment tuning is to leverage the feedback from larger, more powerful LLMs to guide smaller, less capable models. Bai et al. (2022) first propose employing the AI’s feedback to build a harmless AI assistant. They train the reward model using synthetic preference data based on the preference of a pre-trained language model. Building on this, Lee et al. (2023) discover that the RLAIF method can achieve comparable performance with RLHF even when the LLM judge is not strong enough. They also introduced DIRECT-RLAIF, which directly employs an off-the-shelf LLM as the judge model to mitigate reward staleness in reward models. To avoid reward hacking in alignment, Sun et al. (2024a) devise an instructable reward model trained on synthetic preference data. It enables humans to perform RL-time interventions to better align the target policy with human values. Apart from the abovementioned studies, Guo et al. (2024) introduce online AI feedback (OAIF), directly utilizing the preference signals from an annotation model to train the target model. There are also works that utilize multi-agent cooperation for better judgment in alignment tuning. Arif et al. (2024) and Sengupta et al. (2024) construct a synthetic preference optimization dataset using multi-agent workflows and adopt LLMs as judges with diverse prompting strategies and pipelines. Similarly, Li et al. (2024i) employ multiple LLMs to debate with each other, iteratively improving response quality, while creating a judge LLM to select preferred responses for enhanced instruction tuning. To align the generated code with human preference, Weyssow et al. (2024) introduce CodeUltraFeedback, a preference coding dataset constructed using the LLM-as-a-Judge methodology. This synthetic dataset is later used to fine-tune and align small code LLMs using SFT and DPO. Recently, Wang et al. (2024d) propose BPO, using GPT-4 as the judge and constructing synthetic pairwise feedback for knowledge depth and breadth balance in the alignment process.
@@ 5.2.2 Self-Judging
Another line of work aims to utilize the preference signal from the same LLM for self-improving. Yuan et al. (2024c) first propose the concept of a self-rewarding LLM, where pairwise data is constructed by having the LLM itself serve as the judge. Following them, Wu et al. (2024a) introduce meta-rewarding that judges the LLMs’ judge and uses the feedback to refine their judging skills. Their LLM-as-a-meta-judge approach significantly enhances the model’s capacity to evaluate and follow instructions. To improve synthetic data quality, Pace et al. (2024) combine the Best-of-N and Worst-of-N sampling strategies and introduce the West-of-N approach. Lee et al. (2024) devise Judge augmented Supervised Fine-Tuning (JSFT) to train a single model to act as both a policy and a judge. To fully utilize this judge model, they also propose the Self-Rejection by Tournament method to select the best response in the inference time. Unlike the above approaches, which use the LLM as a judge for pairwise data construction, Tong et al. (2024) apply LLM-as-a-judge as a self-filtering method to ensure the quality of synthetic data pairs in alignment tasks for reasoning. To reduce the computation overhead in pairwise judgment, Zhai et al. (2024) propose a ranked pairing method for self-preferring language models, which accelerates the comparison process by measuring the strength of each response against a baseline. Liu et al. (2024c) introduce meta-ranking, enabling weaker LLMs Sun et al. (2024b) to serve as reliable judges and provide trustworthy feedback. They also apply their meta-ranking method in post-SFT training, combining it with Kahneman-Tversky Optimization (KTO) for improved alignment. To enhance the quality of the synthetic instruction tuning data, Liang et al. (2024c) introduce an iterative self-enhancement paradigm (I-SHEEP). During training, they adopt LLM-as-a-judge to score the synthetic responses and set a threshold to collect high-quality query-response pairs for the subsequent training iteration. Recently, Yasunaga et al. (2024) propose to combine LLM-as-a-judge and data synthesis and successfully build human-aligned LLMs with a few annotation.
Several works have also employed LLM-as-a-judge in specific domains or for particular attributes. Zhang et al. (2024h) propose a self-evaluation mechanism, judging responses’ factuality by generating question-answer pairs. They then utilize these self-annotated responses to fine-tune the model via the DPO algorithm for better factuality. In robotics, Zeng et al. (2024) iteratively update the reward function with the self-ranking responses from LLMs, boosting learning efficiency without human supervision. In the multimodal domain, Ahn et al. (2024) propose iterative self-retrospective judgment (i-SRT), which employs self-reflection to improve both response generation and preference modeling.
@@ 5.3.1 Traditional Retrieval
Beyond general retrieval tasks, LLMs have demonstrated their utility as judges in specialized applications. For example, Ma et al. (2024a) outline a few-shot workflow that employs a general-purpose LLM for relevance judgments in legal information retrieval. This model achieves high consistency with expert annotations by breaking down tasks into stages, facilitating the integration of expert reasoning to improve relevance assessment accuracy in legal contexts.
@@ 5.3.2 Retrieval-Augmented Generation (RAG)
RAG systems tailored to specific domains reveal the potential for LLMs to navigate complex queries by integrating specialized knowledge structures. In domain-specific retrieval, Wang et al. (2024b) present BIORAG, an advanced RAG framework that enhances vector retrieval with hierarchical knowledge structures. BIORAG adopts a self-aware evaluated retriever to continuously judge the adequacy and relevance of the information it has collected, thus improving the accuracy of the retrieved document. For biomedical research, Li et al. (2024c) introduce DALK (Dynamic Co-Augmentation of LLMs and Knowledge Graphs), a novel system that combines an LLM with a continuously evolving Alzheimer’s Disease (AD) knowledge graph derived from the scientific literature. Using the novel self-aware knowledge retrieval method, DALK employs the judging capability of LLMs to do noise filtering, enhancing the LLM’s inferential performance in AD-related queries. Similarly, Jeong et al. (2024) propose Self-BioRAG, a framework that adapts RAG principles to biomedical applications. Self-BioRAG employs the LLM to select the best evidence and generate the answer based on the selected evidence and encoded knowledge. Recently, Déjean (2024) distill an LLM-as-a-judge to extract retrieval results from the LLMs’ parametric memory.
@@ 5.4.1 Reasoning Path Selection
Some works focus on sample-level selection in the reasoning process.  Gao et al. (2023a) introduce a strategy evaluator to further evaluate the candidate strategies on a validation set.  Kawabata and Sugawara (2024) introduce REPS (Rationale Enhancement through Pairwise Selection), judging and selecting valid rationales via pairwise self-evaluation with LLMs and training a verifier based on these data. Another reasoning path selection problem for LLMs is diversity, Lahoti et al. (2023) find that LLMs grasp the concept of diversity and can identify aspects where responses lack diversity. By selecting and aggregating multiple critiques, LLMs can achieve similar gains, compared to multiple iterations of critique and revision. In multi-agent collaboration frameworks, Liang et al. (2023) propose multi-agent debating (MAD), a novel paradigm to facilitate debating and discussion among multiple agents. They leverage the judge LLM to select the most reasonable response as the final output at the end of the debating process. Similarly, Li et al. (2024b) propose new roles in layer-based multi-agent collaboration, employing the judge LLM to select the high-quality and reasonable response and thus enhancing the whole system’s token utilizing efficiency by a considerable margin.
@@
Additionally, there are also many works that focus on step-level reasoning path selection, leveraging the judge LLM as the process reward model (PRM) to evaluate state scores. Creswell et al. (2023) decompose the reasoning process into Selection and Inference. In the selection step, they leverage the LLM itself to judge and evaluate each potential reasoning trace, selecting the appropriate one for the following inference step. Xie et al. (2024b) propose the Kwai-STaR framework, which transforms LLMs into state-transition reasoners to judge and select the best reasoning state for themselves in mathematical reasoning. Lightman et al. (2023) train a LLM as the PRM to conduct inference-time supervision and perform best-of-N sampling strategy during reasoning stage. Following them, Setlur et al. (2024) further propose process advantage verifiers (PAVs), generating rewards based on the changes in the likelihood of producing correct responses in the future. Other works simulate advanced cognitive structures as the reasoning process. Hao et al. (2023) employ LLMs as a world model to simulate the state of the environment and perform Monte Carlo Tree Search (MCTS) to improve performance on tasks requiring deliberate path selection. Furthermore, Besta et al. (2024) takes the outputs generated by LLMs as an arbitrary graph. An LLM thought is modeled as a vertex, while an edge is a dependency between thoughts. This framework enables a systematic judge of coherence and logical reasoning for each reasoning state. Yao et al. (2023a) propose Tree-of-Thoughts (ToT), where each thought serves as an intermediate step toward problem-solving. It decomposes reasoning into steps, self-evaluates and judges progress at each state, and uses search algorithms with LMs to judge thought paths through lookahead and backtracking. Besides, there are also studies train critique-based LLM judges Ankner et al. (2024); Yu et al. (2024b); Wang et al. (2024c); Lan et al. ; Xie et al. (2024b); Yao et al. (2024) which provide fine-grained verbal feedback to boost the reasoning process.
@@
EvalBiasBench Park et al. (2024) Alignment 80 Human Accuracy Various Bias
@@
Arena-Hard Auto Li et al. (2024k) Alignment 500 GPT-4-Turbo Separability, Agreement Challenging
@@ 5.4.2 Reasoning with External Tools
Yao et al. (2023b) first propose to use LLMs in an interleaved manner to generate reasoning traces and task-specific actions. Reasoning traces help the model to judge and update action plans, while actions enable it to interact with external sources. Further, Auto-GPT was introduced by Yang et al. (2023) to deliver more accurate information with LLM-as-a-judge for tool-using. By equipping with a range of external complex tools, LLMs become more versatile and capable, improving planning performance by judging and reasoning which tools to use. Sha et al. (2023) explore LLMs’ potential in reasoning and judging, employing them as decision-making components for complex autonomous driving scenarios that require human commonsense understanding. Zhou et al. (2024c) utilize a self-discovery process where LLMs perform judgment based on the given query and select the most feasible reasoning structure for the following inference stage.
@@ 6.1 General Performance
  • JUDGE-BENCH Tan et al. (2024a), DHP Wang et al. (2024g), RewardBench Lambert et al. (2024), SOS-BENCH Penfever et al. (2024) and JuStRank Gera et al. (2024), which operate at larger scales, utilizing metrics like Cohen’s kappa, Discernment Score, and normalized accuracy to benchmark general LLM performance.
@@ 6.2 Bias Quantification
Mitigating bias in LLM judgments is critical to ensuring fairness and reliability Schroeder and Wood-Doughty (2024). Typical benchmarks include EvalBiasBench EvalBiasBench and CALM Ye et al. (2024a), focus explicitly on quantifying biases, including those emerging from alignment and robustness under adversarial conditions. Besides, Shi et al. (2024a) evaluate metrics such as position bias and percent agreement in question-answering tasks.
@@ 6.3 Challenging Task Performance
Benchmarks designed for difficult tasks push the boundaries of LLM evaluation. For example, Arena-Hard Auto Li et al. (2024k) and JudgeBench Tan et al. (2024a) select harder questions based on LLMs’ performance for conversational QA and various reasoning tasks, respectively. CALM Ye et al. (2024a) explores alignment and challenging scenarios, using metrics such as separability, agreement, and hacked accuracy to evaluate performance in manually identified hard data sets.
@@ 6.4 Domain-Specific Performance
Domain-specific benchmarks provide task-focused evaluations to assess LLMs’ effectiveness in specialized contexts. Concretely, Raju et al. (2024) measure separability and agreement across tasks, leveraging metrics such as Brier scores for insights in specific domains such as coding, medical, finance, law and mathematics. CodeJudge-Eval Zhao et al. (2024a) specifically evaluates LLMs for judging code generation with execution-focused metrics such as accuracy and F1 score. This idea has also been adopted by several following works in code summarization and generation evaluation Wu et al. (2024b); Yang et al. (2024); Tong and Zhang (2024); Lin et al. (2024); Gomes et al. (2024); Yu et al. (2024d)
@@ 7.1 Bias & Vulnerability
In addition to these general biases, specific evaluative biases emerge when LLMs act as judges. Order Bias is a prominent issue where the sequence of candidates influences preferences Zheng et al. (2023); Wang et al. (2023c); Koo et al. (2023a); Shi et al. (2024b). This bias can distort evaluation outcomes, particularly in pairwise comparisons, and is more pronounced when the quality gap between competing responses is small Shi et al. (2024b); Wang et al. (2023c). Egocentric Bias arises when LLMs favor outputs generated by the same model, compromising objectivity Liu et al. (2023c); Koo et al. (2023a); Wataoka et al. (2024); Panickssery et al. (2024). This issue is particularly pronounced when evaluation metrics are designed using the same model, leading to inflated scores for outputs from the originating model Liu et al. (2023c). Length Bias, another prevalent challenge, skews evaluations by disproportionately favoring longer or shorter responses regardless of quality Zheng et al. (2023); Koo et al. (2023a). Additional biases, such as Misinformation Oversight Bias, Authority Bias, and Beauty Bias, further complicate LLM evaluations. For instance, misinformation oversight bias reflects a tendency to overlook factual errors, authority bias favors statements from perceived authoritative sources, and beauty bias prioritizes visually appealing content over substantive quality Chen et al. (2024b, d); Stephan et al. (2024). Similarly, Verbosity Bias shows a preference for lengthier explanations, often equating verbosity with quality, which can mislead the judgment process Yuan et al. (2024a). Furthermore, Sentiment Bias skews evaluations based on emotional tone, favoring responses with positive phrasing Ye et al. (2024a).
LLM judges are also highly susceptible to adversarial manipulations. Techniques like JudgeDeceiver highlight the risks posed by optimization-based prompt injection attacks, where carefully crafted adversarial sequences can manipulate LLM judgments to favor specific responses Shi et al. (2024a); BENCHMARK ; Banerjee et al. (2024). Similarly, universal adversarial phrases can drastically inflate scores in absolute scoring paradigms, revealing vulnerabilities in zero-shot assessment setups Liusie et al. (2023); Raina et al. (2024); Doddapaneni et al. (2024). These manipulations raise concerns about the reliability of LLM judges in high-stakes scenarios such as leaderboards and academic or legal assessments Shi et al. (2024a); Raina et al. (2024); Zhang et al. (2024e); Jiang et al. (2024).